OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0 Can i have help its urgent

OpenStudy (zehanz):

Are you sure is is not cos²x-cosx=0?

OpenStudy (zehanz):

(last week I answered the same question, it was cos²x eventually...)

OpenStudy (anonymous):

no it says cos 2x

OpenStudy (zehanz):

If is is cos 2x - cos x = 0, write as cos 2x = cos x. Then remember the general form: \[\cos a = \cos b \Leftrightarrow a=\pm b + 2k \pi\]

OpenStudy (zehanz):

So you would get:\[2x=x + 2k \pi \vee 2x=-x + 2k \pi\] Now, these are two separate equations that can easily be solved. Because the interval is [0, 2pi) you will not have many solutions left.